OpenStudy (4meisu):

Help please: trig! Find exactly the 2 principal angles you would use to find solutions to: 2cos(3(x- π/4 ))=√3

OpenStudy (dumbcow):

start with this \[\cos \theta = \frac{\sqrt{3}}{2}\] which 2 angles solve this?

OpenStudy (4meisu):

30 and 150?

OpenStudy (dumbcow):

correct now sub back in stuff in the cosine \[3(x-45) = 30\] solve for x repeat for the 150 to get 2nd solution

OpenStudy (4meisu):

So I sub in 30 into 3(x-45)=30 to get my first principle angle?

OpenStudy (dumbcow):

no sorry i wasn't clear \[\cos \theta = \frac{\sqrt{3}}{2}\] \[\theta = 30,150\] however from the original equation, \[\theta = 3(x-45)\] therefore \[3(x-45) = 30,150\] solve for "x" .... nothing to sub in

OpenStudy (4meisu):

Aaah, I see. Thank you!